# Fuzzy Information Evolution with Three-Way Decision in Social Network Group Decision-Making

### v1

@@ Abstract
In group decision-making (GDM) scenarios, uncertainty, dynamic social structures, and vague information present major challenges for traditional opinion dynamics models. To address these issues, this study proposes a novel social network group decision-making (SNGDM) framework that integrates three-way decision (3WD) theory, dynamic network reconstruction, and linguistic opinion representation. First, the 3WD mechanism is introduced to explicitly model hesitation and ambiguity in agent judgments, thereby preventing irrational decisions. Second, a connection adjustment rule based on opinion similarity is developed, enabling agents to adaptively update their communication links and better reflect the evolving nature of social relationships. Third, linguistic terms are used to describe agent opinions, allowing the model to handle subjective, vague, or incomplete information more effectively. Finally, an integrated multi-agent decision-making framework is constructed, which simultaneously considers individual uncertainty, opinion evolution, and network dynamics. The proposed model is applied to a multi-UAV cooperative decision-making scenario, where simulation results and consensus analysis demonstrate its effectiveness. Experimental comparisons further verify the advantages of the algorithm in enhancing system stability and representing realistic decision-making behaviors.
@@ I Introduction
In multi-agent decision-making, information is often highly uncertain. Traditional centralized information fusion and decision-making systems often face problems such as isolated information and poor coordination. To improve local assessment capabilities and overall situation awareness, distributed information fusion and decision-making mechanisms based on multi-agent systems have attracted increasing attention in recent years [1].
@@
Opinion dynamics models have been widely used as a fundamental framework for distributed information fusion in multi-agent systems. By simulating how individual agents update their information, these models effectively show collective behavior and the process of consensus formation in uncertain environment [2]. In [3], the DeGroot and Hegselmann–Krause (HK) models were combined to propose a hybrid opinion dynamics (HOD) model in the large-scale group decision-making (LSGDM) framework. In [4], a self-confidence evolution model, which encompassed the self-confidence levels of one’s group mates and the passage of time, was proposed. To consider the trust relationship in social network group decision making (SNGDM), a new discount-based uninorm trust propagation operator with 2-tuple linguistic was presented in [5]. Following the idea of the personalized individual semantics (PIS) model, the PIS-based linguistic opinions dynamics (PIS-LOD) model was proposed within the framework of bounded confidence effects [6]. In [7], a novel opinion dynamics model called interval opinion dynamics with dynamic bounded confidence was proposed, in which agents expressed their opinions in numerical intervals and the bounded confidences varied within a specified range over time. By taking different numerical interval opinions and uncertainty tolerances into account, a numerical interval opinion dynamics model was proposed to investigate the process of forming collective opinions in a group of interacting agents [8]. Inspired by the DeGroot model, a novel group decision-making (GDM) model with opinions evolution was established, and a consensus model with minimum adjustments was provided to obtain the optimal adjusted initial opinions and collective consensus opinion [9]. Besides, an opinion similarity mixed (OSM) model and a structural similarity mixed (SSM) model were proposed in [10], in which the strong and weak relations between individuals were considered. Based on the social network DeGroot (SNDG) model, the authors proved that the speed of consensus reaching was subject to the largest self-confidence level of opinion followers and the speed of consensus reaching was subject to the top two self-confidence levels of opinion leaders [11]. In [12], a conflict resolution method considering the trust propagation, conflict detection, and alternatives selection was proposed.
@@
As shown in the above analysis, opinion dynamics models have been widely applied across various fields. However, most of these models assume that agents must always make definite choices. For example, in the HK model, if the opinion difference between agents exceeds a given threshold, they will not interact or update their information. However, in many cases, the actual opinion difference may be only slightly larger than the threshold, yet it completely blocks information exchange, which is often unrealistic in practical situations. To overcome this limitation, this study incorporates the three-way decision (3WD) theory. 3WD has demonstrated unique advantages in various decision-making environments. In [13], the update problem of 3WD with dynamic variation of scales in incomplete multi-scale information systems was studied, where the updating mechanisms of decision granules induced by the similarity relation were exploited through cut refinement and coarsening via attribute value taxonomies. Considering that the existing 3WD models could not effectively handle incomplete multiattribute decision-making (MADM) problems in real life, the preference of decision-makers for each alternative was considered and the concept of predecisions was introduced, whereby an incomplete fuzzy decision system (IFDS) was obtained in [14]. In [15], a kind of 3WD-MADM with probabilistic dominance relations was proposed, where the two state sets were developed by virtue of probabilistic dominance classes. Also, a probabilistic linguistic 3WD method based on regret theory (RT), namely PL-TWDR, was proposed for MADM problems with probabilistic linguistic term sets (PLTSs), considering the effective rationality of a decision-maker in complex decision environments [16]. In [17], the concepts of positive, neutral, and negative alliances with two thresholds were first provided, and then three-way conflict analysis based on the Bayesian minimum risk theory was conducted, where examples were explored to show how the positive, neutral, and negative alliances were computed with a Pythagorean fuzzy loss function given by an expert.
@@
Moreover, the structure of social network is also essential in GDM. In [18], a cooperation index was formulated to identify the non-cooperative behavior, based on which a non-cooperative behavior transformation method was investigated and a bidirectional feedback mechanism was provided for SNGDM. In [19], trust risks in the conflict-eliminating process (CEP) of SNGDM were addressed through third-party monitoring, and a trust risk analysis-based conflict-eliminating model for SNGDM was developed. In [20], an interaction indicator was exploited to represent peer interaction effort (PIE), individual social cooperation networks (ISCNs) were constructed with the log-sigmoid transition technique, and a novel game-theoretic expert importance evaluation model guided by cooperation effects was finally proposed. To address the challenge that decision-makers in a community setting usually exhibited complex social preferences and intricate social interactions, a minimum cost consensus-based SNGDM approach considering altruism-fairness preferences and ordered trust propagation was designed in [21].
Despite the wide application of opinion dynamics models in distributed decision-making tasks, several key limitations remain. First, many classical models such as the DeGroot and HK frameworks assume that agents must always make definite decisions, typically accepting or rejecting a certain view. This binary assumption overlooks the hesitation and ambiguity that often arise in real-world environments, especially when information is incomplete or conflicting. Although the HK model introduces a confidence threshold to support selective interactions, the results are highly sensitive to small changes in this parameter, which reduces its robustness in dynamic settings. Second, most existing models do not incorporate a mechanism to explicitly handle indecision or uncertainty during the opinion updating process. As a result, agents are often forced to take positions even when the available information does not justify a clear choice. This can lead to premature convergence or extreme polarization. Third, the role of network structure is often simplified or treated as static, whereas in many practical scenarios, communication and influence networks evolve dynamically. Agents do not interact with all others equally; rather, their connections change based on factors such as opinion similarity, trust, or communication cost. Neglecting these structural dynamics limits the model’s ability to reflect realistic social interactions and adaptive coordination.
@@
    Introducing 3WD to describe uncertainty in agent judgments. Unlike traditional opinion dynamics models that use binary decisions, this study applies the 3WD mechanism (“accept,” “reject,” and “defer”) to better handle ambiguous or conflicting information. This helps prevent irrational decisions and premature consensus.
@@
    Combining opinion dynamics with a dynamic social network reconstruction mechanism. A connection adjustment rule based on opinion similarity is proposed, allowing agents to adaptively update their communication links. This reflects the dynamic and selective nature of real-world social interactions.
@@
    Using linguistic information to better match real-world scenarios. Instead of using only numerical values, agent opinions are described with linguistic terms. This improves the model’s ability to deal with vague, subjective, or incomplete information, making the decision-making process more practical to explain.
  4. 4.
    Developing an integrated decision-making framework for multi-agent coordination. The framework considers uncertainty in individual decisions, opinion evolution in groups, and changes in network structure, which improves the system’s performance in complex environments.
@@ II-B Linguistic Scale Functions (LSFs)
[…]
@@ II-C DeGroot Model
Definition 3 [25]: In the DeGroot model, the agents unconditionally trust the other agents. Suppose that $E=\{e_{1},e_{2},\cdots,e_{N}\}$ indicates the agents. $t\in\{1,2,\cdots,T\}$ indicates the discrete variable representing the number of rounds of opinion evolution. For $i\in N$, $y_{i}(t)\in[0,1]$ is the numeric opinion at time $t$. The weight offered by the agent $e_{i}$ to $e_{j}$ is expressed as $\omega_{j}$, where $0\leq\omega_{j}\leq 1$ and $\sum_{j=1}^{N}\omega_{j}=1$ for all $i\in N$. The opinions evolution rule is:
@@ II-D HK Bounded Confidence Model
When the agent $e_{i}$ update the opinion at time $t$, the weight of the agent $e_{j}$ is:
@@
Let $x_{i}(t+1)$ indicates the opinion of $e_{i}$ at time $t+1$, then
@@ II-E Social Network
Definition 5 [28, 29, 30]: A social network is $G(V,E)$. $V=\{v_{t},\;t\in M\}$ expresses vertices. $E=\{e_{ts},\;t,s\in M\}$ indicates the edges. The edge $e_{ts}=(v_{t},v_{s})$ indicates the relationship between $v_{t}$ and $v_{s}$. $F=(f_{ts})_{m\times m}$ can describe the social network.
@@
Definition 7 [33]: The 3WD model includes two states $X=\{C,\neg C\}$ and three actions $H=\{h_{A},h_{D},h_{R}\}$, where $C$ and $\neg C$ represent for a “good” state and a “bad” state, respectively. $h_{A}$, $h_{D}$, and $h_{R}$ represent the three domains in classifying an alternative $a_{i}$, that is, $a_{i}\in POS(C),\,a_{i}\in BOU(C),\,\text{and}\;a_{i}\in NEG(C)$, respectively. Furthermore, $POS(C),\,BOU(C),\,\text{and}\;NEG(C)$ correspond to the positive domain, the boundary domain and the negative domain, respectively. When $a_{i}\in C$, let $\lambda_{AP}$, $\lambda_{DP}$, and $\lambda_{RP}$ denote the losses of taking the actions $h_{A}$, $h_{D}$, and $h_{R}$. Similarly, when $a_{i}\in\neg C$, let $\lambda_{AN}$, $\lambda_{DN}$, and $\lambda_{RN}$ represent the losses of taking the actions $h_{A}$, $h_{D}$, and $h_{R}$. Then, the loss functions of $a_{i}$ are given in Table I, where $C(P)$ and $\neg C(P)$ are represented by $P$ and $N$ standing for $C$ and $\neg C$, respectively. For any alternative $a_{i}\in A$, the equivalence class of $a_{i}$ is represented as $[a_{i}]$. $Pr(C\mid[a_{i}])$ and $Pr(\neg C\mid[a_{i}])$ represent the conditional probability of the alternative $a_{i}$ belongs to and not belonging to $C$, respectively. Then, the expected losses $R(h_{I}\mid[a_{i}])$ for the three actions ($I=A;D;R$) are:
@@ III Linguistic 3WD Opinion Dynamics Model over
To better model the interaction between opinion evolution and network structure in GDM, this paper proposes a co-evolutionary model that combines the 3WD mechanism with a dynamic network update process. Traditional models often use a binary decision rule, where agents either fully accept or completely reject neighbors based on a fixed confidence bound. However, such a hard-boundary approach cannot fully reflect the diversity of real-world social behavior. In this work, we introduce the 3WD mechanism, which divides the opinion distance between two agents into three regions: if the distance is less than a strict acceptance threshold $\alpha$, the neighbor is deterministically accepted; if the distance exceeds a rejection threshold $\beta$, the neighbor is rejected; and if the distance falls between $\alpha$ and $\beta$, the agent accepts the neighbor with a certain probability. This probability is controlled by an exponential decay function $p=\exp(-\lambda(d_{ij}-\alpha))$, where $d_{ij}$ is the opinion distance and $\lambda$ is a decay factor. This design reflects the psychological tendency that people are more likely to accept opinions that are closer to their own, while opinions farther away are less likely to be accepted. The 3WD mechanism allows the model to capture hesitation, selective trust, and partial agreement, which are common in real social interactions. In addition, to break the limitation of a static network, we introduce a dynamic update mechanism based on cognitive similarity. This mechanism allows the connections between agents to change over time, depending on their current opinions. Specifically, if two agents are not connected but their opinion distance is smaller than a threshold $\delta_{\text{add}}$, a link is added between them with probability $p_{\text{add}}$. On the other hand, if two connected agents show a large opinion difference that exceeds $\delta_{\text{cut}}$, their link may be removed with probability $p_{\text{cut}}$. In this way, the network structure evolves according to the cognitive relationships among agents.
@@
  • $A_{ij}(t)$: Social adjacency matrix ($1$ if connected, 00 otherwise)

### v2

@@ Abstract
In group decision-making (GDM) scenarios, uncertainty, dynamic social structures, and vague information present challenges for traditional opinion dynamics models. To address these issues, this study proposes a novel social network group decision-making (SNGDM) framework that integrates three-way decision (3WD) theory, dynamic network reconstruction, and linguistic opinion representation. First, the 3WD mechanism is introduced to explicitly model hesitation and ambiguity in agent judgments, thereby preventing irrational decisions. Second, a connection adjustment rule based on opinion similarity is developed, enabling agents to adaptively update their communication links and better reflect the evolving nature of social relationships. Third, linguistic terms are used to describe agent opinions, allowing the model to handle vague and incomplete information more effectively. Finally, an integrated multi-agent decision-making framework is constructed, which simultaneously considers individual uncertainty, opinion evolution, and network dynamics. The proposed model is applied to a multi-UAV cooperative decision-making scenario, where simulation results and consensus analysis demonstrate its effectiveness. Experimental comparisons further verify the algorithm’s advantages in enhancing system stability and representing realistic decision-making behaviors.
@@ I Introduction
In multi-agent decision-making, information is often highly uncertain. Traditional centralized information fusion and decision-making systems often face problems such as isolated information and poor coordination. To enhance local assessment capabilities and overall situational awareness, distributed information fusion and decision-making mechanisms based on multi-agent systems have garnered increasing attention in recent years [1].
@@
Opinion dynamics models have been widely used as a fundamental framework for distributed information fusion in multi-agent systems. By simulating how individual agents update their information, these models effectively show collective behavior and the process of consensus formation in an uncertain environment [2]. In [3], the DeGroot and Hegselmann–Krause (HK) models were combined to propose a hybrid opinion dynamics (HOD) model in the large-scale group decision-making (LSGDM) framework. In [4], a self-confidence evolution model, which encompassed the self-confidence levels of one’s group mates and the passage of time, was proposed. To consider the trust relationship in social network group decision making (SNGDM), a new discount-based uninorm trust propagation operator with 2-tuple linguistic was presented in [5]. Following the idea of the personalized individual semantics (PIS) model, the PIS-based linguistic opinions dynamics (PIS-LOD) model was proposed within the framework of bounded confidence effects [6]. In [7], a novel opinion dynamics model called interval opinion dynamics with dynamic bounded confidence was proposed, in which agents expressed their opinions in numerical intervals. By taking different numerical interval opinions and uncertainty tolerances into account, a numerical interval opinion dynamics model was proposed to investigate the process of forming collective opinions in a group of interacting agents [8]. Inspired by the DeGroot model, a novel group decision-making (GDM) model with opinion evolution was established, and a consensus model with minimum adjustments was provided to obtain the optimal adjusted initial opinions and collective consensus opinion [9]. Besides, an opinion similarity mixed (OSM) model and a structural similarity mixed (SSM) model were proposed in [10], in which the strong and weak relations between individuals were considered. Based on the social network DeGroot model, the authors proved that the speed of consensus reaching was subject to the most significant self-confidence level of opinion followers and the speed of consensus reaching was subject to the top two self-confidence levels of opinion leaders [11]. In [12], a conflict resolution method considering the trust propagation, conflict detection, and alternative selection was proposed.
@@
As shown in the above analysis, opinion dynamics models have been widely applied across various fields. However, most of these models assume that agents must always make definite choices. For example, in the HK model, if the opinion difference between agents exceeds a given threshold, they will not interact or update their information. However, in many cases, the actual opinion difference may be only slightly larger than the threshold, yet it completely blocks information exchange, which is often unrealistic in practical situations. To overcome this limitation, this study incorporates the three-way decision (3WD) theory. 3WD has demonstrated unique advantages in various decision-making environments. In [13], the update problem of 3WD with dynamic variation of scales in incomplete multi-scale information systems was studied, where the updating mechanisms of decision granules induced by the similarity relation were exploited through cut refinement and coarsening via attribute value taxonomies. Considering that the existing 3WD models could not effectively handle incomplete multiattribute decision-making (MADM) problems in real life, the preference of decision-makers for each alternative was considered, and the concept of predecisions was introduced, whereby an incomplete fuzzy decision system (IFDS) was obtained in [14]. In [15], a kind of 3WD-MADM with probabilistic dominance relations was proposed, where the two state sets were developed by virtue of probabilistic dominance classes. Also, a probabilistic linguistic 3WD method based on regret theory (RT), namely PL-TWDR, was proposed for MADM problems with probabilistic linguistic term sets (PLTSs), considering the practical rationality of a decision-maker in complex decision environments [16]. In [17], the concepts of positive, neutral, and negative alliances with two thresholds were first provided. Then, a three-way conflict analysis based on the Bayesian minimum risk theory was conducted, where examples were explored to show how the positive, neutral, and negative alliances were computed with a Pythagorean fuzzy loss function given by an expert.
@@
Moreover, the structure of the social network is also essential in GDM. In [18], a cooperation index was formulated to identify the non-cooperative behavior, based on which a non-cooperative behavior transformation method was investigated, and a bidirectional feedback mechanism was provided for SNGDM. In [19], trust risks in the conflict-eliminating process (CEP) of SNGDM were addressed through third-party monitoring, and a trust risk analysis-based conflict-eliminating model for SNGDM was developed. In [20], an interaction indicator was exploited to represent peer interaction effort (PIE), individual social cooperation networks (ISCNs) were constructed with the log-sigmoid transition technique, and a novel game-theoretic expert importance evaluation model guided by cooperation effects was finally proposed. To address the challenge that decision-makers in a community setting typically exhibit complex social preferences and intricate social interactions, a minimum cost consensus-based SNGDM approach was designed, considering altruism-fairness preferences and ordered trust propagation [21].
Despite the wide use of opinion dynamics models in distributed decision-making, several important challenges remain. First, some classical models, such as the DeGroot and HK frameworks, assume that agents must always make a clear decision, usually to accept or reject a particular opinion, and do not provide a mechanism to express hesitation or uncertainty explicitly. This forces agents to take a position even when the available information is not enough, which may lead to premature consensus or extreme polarization. For example, in a public policy vote, if a person does not have enough information about the pros and cons of a proposal, they may prefer to stay undecided until more evidence is available. However, in many classical models, this person is still forced to move toward “support” or “oppose” in each update step. The HK model partly addresses this problem by introducing a confidence threshold to limit interactions, but its results are susceptible to small changes in this threshold. For example, when the threshold is set to 0.3, if the opinion difference between two agents is 0.29, they will interact; but if the difference is 0.31, they will refuse to interact. In reality, these two situations are almost the same, but the model produces completely different results. This “boundary effect” renders the model highly sensitive to threshold settings, which reduces its robustness in dynamic environments. Second, some studies treat the network structure as fixed or oversimplified. In reality, communication and influence networks change over time. Agents interact more frequently with some individuals than others, and these connections evolve in response to factors such as opinion similarity, trust, or communication costs. For example, on social media platforms, people may increase interaction if their opinions become closer, or reduce contact and even unfollow, if their views diverge significantly. Models that maintain a static network cannot capture this adaptive connection change, which limits their ability to describe fundamental social interactions and coordination.
@@
    Three-way decision (3WD) mechanism for hesitation and uncertainty. We design an opinion updating rule based on the three-way decision framework, where each agent can choose “accept,” “reject,” or “defer” in response to incoming information. This differs from classical models such as DeGroot or HK that force agents into binary decisions. The mechanism explicitly preserves hesitation intervals, reducing the risk of premature consensus and improving decision robustness under conflicting or incomplete information.
@@
    Dynamic social network reconstruction driven by opinion similarity. We introduce a probabilistic link creation and deletion rule that updates the social network structure according to opinion distances between agents. This adaptive adjustment enables agents to strengthen ties with similar peers and weaken ties with divergent ones, capturing the evolving nature of real-world communication patterns.
@@
    Linguistic opinion representation for vague and subjective information. We represent opinions using linguistic terms mapped from numerical values, instead of using only numerical states. This makes the model better suited for situations where information is subjective, vague, or incomplete. The approach improves interpretability and aligns simulation outputs more closely with real-world decision-making contexts, as validated in multi-agent test scenarios.
@@ II-B Linguistic Scale Functions (LSFs)
Definition 2 [24]: For the linguistic term $h_{j}$ in $H=\{h_{i}|i=0,1,\cdots,2\Phi,|\Phi\in N^{*}\}$, $h_{j}$ represents a possible value for a linguistic variable. If $\theta\in[0,\,1]$ is a numerical value, then the function $f:h_{j}\rightarrow\theta_{j}$, which conducts the mapping from $h_{j}$ to $\theta_{j}\;(j=0,\;1,\;\cdots,\;2t)$, is defined as follows:
@@ II-C DeGroot Model
Definition 3 [25]: In the DeGroot model, the agents unconditionally trust each other. Suppose that $E=\{e_{1},e_{2},\cdots,e_{N}\}$ indicates the agents. $t\in\{1,2,\cdots,T\}$ indicates the discrete variable representing the number of rounds of opinion evolution. For $i\in N$, $y_{i}(t)\in[0,1]$ is the numeric opinion at time $t$. The weight offered by the agent $e_{i}$ to $e_{j}$ is expressed as $\omega_{j}$, where $0\leq\omega_{j}\leq 1$ and $\sum_{j=1}^{N}\omega_{j}=1$ for all $i\in N$. The opinion evolution rule is:
@@ II-D HK Bounded Confidence Model
When the agent $e_{i}$ updates the opinion at time $t$, the weight of the agent $e_{j}$ is:
@@
Let $x_{i}(t+1)$ indicate the opinion of $e_{i}$ at time $t+1$, then
@@ II-E Social Network
Definition 5 [28, 29, 30]: A social network is $G(V,E)$. $V=\{v_{t},\;t\in M\}$ expresses vertices. $E=\{e_{ts},\;t,s\in M\}$ indicates the edges. The edge $e_{ts}=(v_{t},v_{s})$ indicates the relationship between $v_{t}$ and $v_{s}$. $F=(f_{ts})_{m\times m}$ describes the social network.
@@
Definition 7 [33]: The 3WD model includes two states $X=\{C,\neg C\}$ and three actions $H=\{h_{A},h_{D},h_{R}\}$, where $C$ and $\neg C$ represent for a “good” state and a “bad” state, respectively. $h_{A}$, $h_{D}$, and $h_{R}$ represent the three domains in classifying an alternative $a_{i}$, that is, $a_{i}\in POS(C),\,a_{i}\in BOU(C),\,\text{and}\;a_{i}\in NEG(C)$, respectively. Furthermore, $POS(C),\,BOU(C),\,\text{and}\;NEG(C)$ correspond to the positive domain, the boundary domain, and the negative domain, respectively. When $a_{i}\in C$, let $\lambda_{AP}$, $\lambda_{DP}$, and $\lambda_{RP}$ denote the losses of taking the actions $h_{A}$, $h_{D}$, and $h_{R}$. Similarly, when $a_{i}\in\neg C$, let $\lambda_{AN}$, $\lambda_{DN}$, and $\lambda_{RN}$ represent the losses of taking the actions $h_{A}$, $h_{D}$, and $h_{R}$. Then, the loss functions of $a_{i}$ are given in Table I, where $C(P)$ and $\neg C(P)$ are represented by $P$ and $N$ standing for $C$ and $\neg C$, respectively. For any alternative $a_{i}\in A$, the equivalence class of $a_{i}$ is represented as $a_{i}]$. $Pr(C\mid[a_{i}])$ and $Pr(\neg C\mid[a_{i}])$ represent the conditional probability of the alternative $a_{i}$ belongs to and not belonging to $C$, respectively. Then, the expected losses $R(h_{I}\mid[a_{i}])$ for the three actions ($I=A;D;R$) are:
@@ III Linguistic 3WD Opinion Dynamics Model over
To better model the interaction between opinion evolution and network structure in GDM, this paper proposes a co-evolutionary model that combines the 3WD mechanism with a dynamic network update process. Traditional models often use a binary decision rule, where agents either fully accept or completely reject neighbors based on a fixed confidence bound. However, such a hard-boundary approach cannot fully reflect the diversity of real-world social behavior. In this work, we introduce the 3WD mechanism, which divides the opinion distance between two agents into three regions: if the distance is less than a strict acceptance threshold $\alpha$, the neighbor is deterministically accepted; if the distance exceeds a rejection threshold $\beta$, the neighbor is rejected; and if the distance falls between $\alpha$ and $\beta$, the agent accepts the neighbor with a certain probability. This probability is controlled by an exponential decay function $p=\exp(-\lambda(d_{ij}-\alpha))$, where $d_{ij}$ is the opinion distance and $\lambda$ is a decay factor. This design reflects the psychological tendency that people are more likely to accept opinions that are closer to their own, while opinions farther away are less likely to be accepted. The 3WD mechanism allows the model to capture hesitation, selective trust, and partial agreement, which are common in fundamental social interactions. In addition, to break the limitation of a static network, we introduce a dynamic update mechanism based on cognitive similarity. This mechanism allows the connections between agents to change over time, depending on their current opinions. Specifically, if two agents are not connected but their opinion distance is smaller than a threshold $\delta_{\text{add}}$, a link is added between them with probability $p_{\text{add}}$. On the other hand, if two connected agents show a significant opinion difference that exceeds $\delta_{\text{cut}}$, their link may be removed with probability $p_{\text{cut}}$. In this way, the network structure evolves according to the cognitive relationships among agents.
@@
  • $A_{ij}(t)$: Social adjacency matrix ($1$ if connected, 0 otherwise)